OpenStudy (anonymous):

what 2 things affect gravity?

OpenStudy (abhisar):

Hey there! Welcome to OpenStudy :) The force of gravity can be expressed as \(\boxed{\sf F_g=\Large\frac{GM_1M_2}{r^2}}\). Where, \(\sf M_1,M_2~and~r\) are mass of object 1, mass of object 2 and distance between the two objects respectively. Therefore we can say that the two things which affect the force of gravity is \(\bf Mass\) and \(\bf Distance\) between the two objects.

TheSmartOne (thesmartone):

And also \(\large G= 6.67384\times 10^{-11} m^3 \times kg^{-1} s^{-2}\)